OpenStudy (anonymous):

3. Which of the following is a solution to the equation 16= 4x-4? a.-5 b.-4 c.5 d.16 please help. I'll give a medal + fan I wanna know how

OpenStudy (texaschic101):

16 = 4x - 4 --- we want to get the 4x by itself, so add 4 to both sides 16 + 4 = 4x - 4 + 4 --- combine terms 20 = 4x -- now we want x by itself, so we divide both sides by 4 20/4 = 4x/4 5 = x you can even check your answer by subbing it into the problem 16 = 4(5) - 4 16 = 20 - 4 16 = 16 (correct) any questions ?

OpenStudy (anonymous):

@Mimi_Styles its 5

OpenStudy (anonymous):

hmm, I kinda understand it now. and yes I have a few more questions :)

OpenStudy (texaschic101):

ask away :)

OpenStudy (anonymous):

Are you talking to me? Thats like 8th grade math dude..hahah

OpenStudy (anonymous):

4. An art student wants to make a model of a rectangular classroom. The length of the classroom is 2.4 times its width. The length of the student's model is 42 in. What should the width of the model be? a. 17.5in b. 20.5 in c. 83. 6in d. 100.8in

OpenStudy (anonymous):

and um.. sorry if I didn't get it. I never really understood math like that. I was always horrible at it...

OpenStudy (anonymous):

._.

OpenStudy (texaschic101):

L = 2.4W L = 42 42 = 2.4W -- divide by 2.4 42/2.4 = W 17.5 = W

OpenStudy (anonymous):

ohhhhhh I see now ^_^

OpenStudy (texaschic101):

good to hear :) Sometimes these problems can be confusing and it helps to have someone walk you though them. Do you have any more questions ?

OpenStudy (anonymous):

yes I do. if that's okay with you haha

OpenStudy (anonymous):

1. Which of the following is an example of an open equation? a. 5x+7=12 b. 12+(-8)= 4 c.25= 5(5) d. 8+0=0 (I kinda forgotten what a open equation was... oops :(..I feel so dumb)

OpenStudy (texaschic101):

an open equation, if I am not mistaken, is one where there is an unsolved variable. Once you sub in the number for the variable, it is no longer open. Your open equation is : 5x + 7 = 12 because we do not know what x is.

OpenStudy (anonymous):

ohhh, okay. may I ask more questions?

OpenStudy (texaschic101):

you may....lol

OpenStudy (anonymous):

yay! *does a happy dance* thankies :)

OpenStudy (anonymous):

2. Which of the following equation has 3 as a solution? a. -2b= 6 b. 4a- 5= 12 c. 5y= y+8 d. z- 15= -12 :D

OpenStudy (texaschic101):

ask when ready

OpenStudy (anonymous):

Its D

OpenStudy (anonymous):

but why ? like why would it be D?

OpenStudy (anonymous):

Because you can just substitute 3 in for z Then it would be 3-15=-12 And 3-15=-12. What grade are you in?

OpenStudy (anonymous):

>_> <_< -hides in a dark hole-

OpenStudy (anonymous):

I was just asking... ._.

OpenStudy (anonymous):

OMG..I am soooooo Sorry if that sounded rude..I wasn't going for the rude sound!!

OpenStudy (anonymous):

And are you the one with the blonde hair in your profile picture?